There are two machines that produce aluminum cans. The newer machine can produce 5400 cans in 180 minutes. It takes the older machine 270 minutes to produce that many cans. If the two machines work together, how long will it take them to produce 5400 cans?
Answer by mananth(16946)

There are two machines that produce aluminum cans. The newer machine can produce 5400 cans in 180 minutes. It takes the older machine 270 minutes to produce that many cans. If the two machines work together, how long will it take them to produce 5400 cans?
new machine --------- 5400 cans in 3 hours
in one hour this machine will produce 5400/3 = 1800 cans
old machine ----------5400 cans in 4.5 hours
this will produce 1200 cans
together they will produce 1800 + 1200 in one hour = 3000 cans.
3000 cans in 1 hour
5400 cans in ---------------- 5400 / 3000 hours = 1.8 hours ie: 108 minutes
0.8 hours = 0.8 * 60 minutes = 48 minutes so 1.8 hours = 60 +48 minutes = 108 minutes